A tree casts a 25-foot shadow on a sunny day. If the angle of elevation from the tip of the shadow to the top of the tree is 32 degrees, what is the height of the tree to the nearest tenth of a foot?
Found 2 solutions by jim_thompson5910, ewatrrr:
Answer by jim_thompson5910(35256)

tan(angle) = opposite/adjacent

tan(32) = x/25

25*tan(32) = x

x = 25*tan(32)

x = 25*0.624869351909327

x = 15.6217337977332

So the tree is roughly 15.6 feet tall.

Answer by ewatrrr(24785)

 
Hi,
A tree casts a 25-foot shadow on a sunny day.
If the angle of elevation from the tip of the shadow to the top of the tree is 32 degrees,
what is the height of the tree to the nearest tenth of a foot?
opp/adj = h/25ft = tan(32°)
h = tan(32°)(25) = 15.6 ft